OpenStudy (anonymous):

Not sure if I asked alady or not,but can someone please explain how to simplify radicals? I am still lost and my teacher sux..an example off my HW is "r5 x r15"

OpenStudy (anonymous):

r^15 * r^5 ?

OpenStudy (anonymous):

no,the r is supposed to stand for the radical sign

OpenStudy (anonymous):

can you draw it with the draw button down below?

OpenStudy (anonymous):

\[\sqrt{a}*\sqrt{b} = \sqrt{ab}\] \[\sqrt{36} = 6\] \[\sqrt{24} = \sqrt{4}*\sqrt{6} = 2\sqrt{6}\] \[\sqrt{18}*\sqrt{10}=\sqrt{9}*\sqrt{2}*\sqrt{2}*\sqrt{5} = 3*2*\sqrt{5} = 6\sqrt{5}\]Are some examples of the basic tools in the toolchest.

OpenStudy (anonymous):

Thanx! And I tried to post it with the drawing but it wont let me

OpenStudy (anonymous):

Try using the symbols \ [ \sqrt{25} = 5 \ ] only jammed together with no whitespace: \[\sqrt{25} =5 \]